OpenStudy (anonymous):

Complete the identity. sin(α + β)/cosαcosβ = ? A. cot α + cot β B. tan β + tan α C. tan α + tan β D. -tan α + cot β

OpenStudy (anonymous):

Well \[\sin(\alpha + \beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\] Dividing that gives: \[\tan(\alpha)+\tan(\beta)\]

OpenStudy (anonymous):

so it would be c?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thank you